OpenStudy (anonymous):

An aircraft travels with the wind for 120 miles in 0.75 of an hour. The return trip is flown against the wind and takes exactly 1 hour. Which system of linear equations represents x, the speed of the plane in miles per hour, and y, the speed of the wind in miles per hour? Recall the formula d = rt.

OpenStudy (anonymous):

0.75(x + y) = 120 x − y = 120 0.75(x − y)= 120 x + y = 120 120(x + y) = 0.75 120(x − y) = 1 120(x − y) = 0.75 120(x + y) = 1

OpenStudy (anonymous):

boxes are -

OpenStudy (anonymous):

If you think about what these numbers represent than you can see which one is correct. x and y is speed: miles per hour 120 is distance so miles For example you can see that x − y = 120 is nonsense miles/hour - miles/hour = miles

OpenStudy (anonymous):

Oh and 0.75 and 1 is hour

OpenStudy (anonymous):

So the only two equations that can be true are 3 and 4. Just think about what they mean and you will be able to solve it